Find the fourth term of the binomial expansion (x-2y2)5

1
Expert's answer
2021-11-30T18:02:44-0500

"=\frac{5}{0!(5-0)!}x^5(-2y^2)^0+\frac{5}{1!(5-1)!}x^4(-2y^2)^1+\frac{5}{2!(5-2)!}x^3(-2y^2)^2+\frac{5}{3!(5-3)!}x^2(-2y^2)^3\\"



"=x^5-10x^4y^2+40x^3y^4-80x^2y^6"